The answer to each of the questions is a single digit integer ranging from 0 to 9.


 


A gas bubble from an explosion under water oscillates with a time t, which is directly proportional to PadbEc where P is the pressure, d is the density and E is the energy of explosion. Find the value of (a+b)/c.


 

Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

3

 

        We have t = kPadbEc;

 

        \ [t] = [P]a[d]b[E]c

 

        Þ M0L0T1 = [MaL–aT–2a][MbL–3b][McL2cT–2c]

 

                        = Ma+b+cL–a–3b+2cT–2a–3b–2c;

 

        \ a+b+c = 0; –a–2b+2c = 0 & –2a–3b–2c = 1, On solving the above

 

        We get, a = 4/3, b = –1, c = –1/3

 

        \